OpenStudy (just_one_last_goodbye):

Evaluate the following expression using values given:

OpenStudy (anonymous):

Hello :D I actually know how to do this

OpenStudy (just_one_last_goodbye):

lol KAYLA!!!!!

OpenStudy (anonymous):

Lol hola Kevin

OpenStudy (anonymous):

So you have to substitute every x with a 3, every y with a 5 and every z with a -3 Sooo.... hold on

OpenStudy (just_one_last_goodbye):

\[x^3 - 2y^2 - 3x^3 + z^4\]

OpenStudy (just_one_last_goodbye):

same thing just wanted to make it look more cleaner

OpenStudy (just_one_last_goodbye):

if x = 3, y = 5, and z =-3

OpenStudy (anonymous):

\[(3)^{3}-2(5)^{2}-3(3)^{3}+(-3)^{4}\] Easier to solve once it's all numbers

OpenStudy (just_one_last_goodbye):

lol ok ^_^

OpenStudy (anonymous):

So what's 3 * 3 * 3?

OpenStudy (just_one_last_goodbye):

27? we are talking about 3 to the power of 3 right?

OpenStudy (acxbox22):

yes

OpenStudy (anonymous):

Yea lol This will take too long I'll just do it (3)^3 ... 3 * 3 * 3 = 27 -2(5)^2 ... -2 * 5 = -10, -10 * -10 = 100 -3(3)^3 ... -3 * 3 = -9, 9 * 9 * 9 = 729 (-3)^4 ... -3 * -3 * -3 * -3 = 81 Shhh let's pretend I'm doing this right Now add: 27 + 100 + 729 + 81

OpenStudy (acxbox22):

no @kayla_renae you messed up in the second step

OpenStudy (acxbox22):

you have to square 5 first then multiply it by -2

OpenStudy (anonymous):

Ooooh :/

OpenStudy (acxbox22):

so 27-50-3(27)+81 27-50-81+81=27-50=-23

OpenStudy (nubeer):

messed in third stepp too

OpenStudy (anonymous):

So 5 * 5 = 25, 25 * -2 = -50 Yea I see 3 * 3 * 3 = 27, 27 * -3 = -81

OpenStudy (anonymous):

Kevin am I confusing you?? xD I suck at this 27 - 50 - 81 + 81 = ...

OpenStudy (just_one_last_goodbye):

oh lol sorry i was in chat

OpenStudy (anonymous):

Was that question multiple choice?

OpenStudy (acxbox22):

i already showed the steps in my post above

OpenStudy (anonymous):

Yup good job @acxbox22 and ty for correcting me I've been doing it wrong for a while lol

OpenStudy (acxbox22):

np ;)